# Safety through feedback in Constrained RL

### v1

@@ 4.4 Policy Optimization
Post the data collection round, where we sample trajectories $\{\tau\}$ and their corresponding rewards $\{r(\tau)\}$ using $\pi$, we estimate their costs $c(\tau)$ using the inferred cost function. Our proposed method allows for the policy to be updated utilizing any on-policy constrained RL algorithm. In this study, we employ the PPO-Lagrangian algorithm[3] that combines the PPO algorithm [24] with a lagrangian multiplier to ensure safety.
@@
1:Input: cost threshold $c_{0}$, segment length $k$, novelty criterion $e$
@@
11:              for each segment $\tau_{j:j+k}\in\tau^{i}$ do
12:                  Obtain feedback $y^{safe}$ for the segment $\tau^{i}_{j:j+k}$.
13:                  $D\leftarrow D\cup\{((s,a),y^{safe})\ \hskip 5.0pt\forall(s,a)\in\tau^{i}_{j:j+k}\}$
14:                  $d(\phi(s))\leftarrow d(\phi(s))+1\hskip 5.0pt\forall s\in\tau^{i}_{j:j+k}$triangleright Update the densities
@@
22:     Infer costs for $\{c_{\theta}(\tau_{i})\}$ for all $\tau^{i}\sim\{\tau\}$.
@@ 5 Experiments
We investigate the following questions in our experiments: 1) Does RLSF succeed in effectively learning safe behaviours? 2) Can the inferred cost function be transferred across agents in the same task? and 3) How does the proposed novelty based sampling scheme compare with other methods used in the literature?
@@ 5.1 Experiment Setup
Table 1: Performance of different algorithms on the Safety Benchmarks. The first $7$ environments represent the hard constraint case. The remaining environments illustrate the soft constraint case, with values in brackets indicating the cost threshold. Each algorithm is run for $6$ independent seeds.
@@
Point Circle Return $45.26$ $46.09$ $36.20\pm 3.95$ $22.26\pm 9.59$ $36.42\pm 1.78$
@@
Car Circle Return $14.34$ $15.21$ $5.18\pm 2.48$ $6.34\pm 2.87$ $9.37\pm 0.97$
@@
Biased Pendulum Return $717.43$ $983.27$ $495.58\pm 160.84$ $577.15\pm 184.31$ $721.48\pm 111.49$
@@
Blocked Swimmer Return $22.62$ $21.05$ $86.96\pm 10.69$ $2.15\pm 8.58$ $16.09\pm 1.44$
@@
HalfCheetah Return $2786.71$ $2497.82$ $3031.7\pm 336.48$ $257.34\pm 147.35$ $2112.63\pm 161.26$
@@
Hopper Return $1705.00$ $1555.25$ $1097.57\pm 56.35$ $990.08\pm 8.66$ $1408.71\pm 27.3$
@@
Walker2d Return $2947.25$ $2925.23$ $2195.94\pm 134.21$ $993.38\pm 17.69$ $2783.29\pm 57.51$
@@
Point Goal Return $26.16$ $26.1$ $1.61\pm 1.8149$ $10.86\pm 4.1$ $24.65\pm 0.59$
Cost $(40.0)$ $34.19$ $31.83$ $30.57\pm 13.29$ $52.76\pm 12.85$ $35.08\pm 1.08$
Car Goal Return $27.37$ $26.44$ $1.05\pm 2.83$ $10.88\pm 7.1$ $24.28\pm 2.1$
Cost $(40.0)$ $41.67$ $35.41$ $34.71\pm 9.87$ $33.33\pm 11.26$ $41.25\pm 2.27$
Point Push Return $6.00$ $10.84$ $0.16\pm 0.14$ $3.63\pm 1.77$ $2.68\pm 1.03$
Cost $(35.0)$ $26.08$ $26.96$ $22.89\pm 5.95$ $45.43\pm 3.86$ $30.51\pm 3.4$
Car Push Return $3.07$ $2.68$ $-3.04\pm 3.3$ $1.56\pm 0.46$ $1.54\pm 0.51$
Cost $(35.0)$ $20.53$ $20.95$ $23.25\pm 7.78$ $36.55\pm 1.48$ 27.69$\pm$ 1.19
@@
We evaluate RLSF on multiple continuous control benchmarks in the Safety Gymnasium environment [15] and Mujoco simulator based environments introduced in [28]. The Circle, Blocked Swimmer and Biased Pendulum environments constrain the position of the agent whereas the Half Cheetah, Hopper and Walker-2d environments constrain the velocity of the agent. The Goal and Push tasks which are the most challenging contain static and dynamic obstacles that the agent must avoid while completing the task. These environments represent safety challenges that an agent may encounter in the real world. Additionally, experiment on the Driver simulator introduced in [19] which introduces two scenarios that an autonomous driving agent is likely to encounter on the highway- the lane change and a blocked path. Here, the cost function considers multiple variables such as speed, position and distance to other cars. We also introduce a third scenario- overtaking in a two lane highway. A detailed description of the tasks can be found in the Appendix B.
@@
We split the environments into two cases: (1) hard constraint case ($c_{max}=0$) where the safety of the policy is measured in terms of the cost violation (CV) rate defined as the average number of cost violations per state and (2) soft constraint case where $c_{max}>0$. We use an automated script that uses the underlying cost function to simulate the feedback received from the evaluator.
@@
We compare the performance of our algorithm against the following baselines: Self Imitation Safe Reinforcement Learning (SIM) [14]: SIM is a state-of-the-art method in constrained RL. It also supports the case where feedback is elicited from an external evaluator. Similar to RLSF, the method consists of two stages, a data collection/feedback stage and a policy optimization stage. In the first stage, a trajectory $\tau$ is labelled as good if $[r(\tau)\geq r_{good}\hskip 5.0pt\cap\hskip 5.0pt\mathbb{I}[c(\tau)\leq c_{max}]$, and labelled as bad if $[r(\tau)\geq r_{bad}\hskip 5.0pt\cup\hskip 5.0pt\mathbb{I}[c(\tau)\geq c_{max}]$, where $r_{good}$ and $r_{bad}$ are predefined thresholds on the reward. The information $\mathbb{I}[c(\tau)\leq c_{max}]$ is received from feedback. The idea is then to imitate the good trajectories and stay away from the bad trajectories. If $\rho^{\pi}$ is the occupancy measure of the current policy $\pi$ and $\rho^{G},\rho^{B}$ are the occupancy measures of the good and bad trajectories respectively, then, $\pi$ is optimized as,
@@
As an upper bound, we compare the performance of our algorithm to the case where the cost function is known: PPO-Lagrangian (PPOLag) [3] and SIM with known costs (SIMKC) [14]. The goal here is not to beat their performance but to match them, hence when reporting the results of these algorithms we report the best performing seed across the different seeds.
All results presented use novelty-based sampling unless stated otherwise. We use a segment length of $1$ in the Driver, Goal and Push environments. This is because the Safety Goal and Push environments contain small obstacles that the agent interacts with for very brief periods of time, hence requiring more fine-grained feedback. An example of this is present in Figure 8 in the Appendix. In the Driver environments, a randomly initialized policy was highly unsafe. Thus a long segment length would force the evaluator to label every segment unsafe, making cost inference infeasible. In all other environments, the segment length is equal to the length of the episode. Details on the number of queries generated for feedback can be found in Table 4 in the Appendix. We provide the two baseline methods (with unknown costs) with an advantage, whereby feedback is received for every trajectory generated.
@@ 5.2 Cost Inference across various tasks
Table 1 contains the performance of the different algorithms on the benchmark environments. RLSF significantly outperforms the two baselines in terms of reward and safety in all of the environments 222except the Car Goal environment where RLSF and PPOLag marginally exceed the threshold. RLSF comes to within $\sim 80\%$ of the performance of the best run of PPOLag in $7/11$ environments thereby underscoring its effectiveness in learning safe policies.
Figure 1: Cost Violation rate of different algorithms in driving environment. Each algorithm is run for $6$ independent seeds, curves represent the mean and shaded regions represent the standard error.
@@
We demonstrate the applicability of our method in learning safe autonomous driving policies. We compare the performance of our algorithm against two baselines: a naive PPO policy that solely maximizes reward and the PPO-Lag algorithm. Figure 1 presents these results. Notably, our method exhibits comparable safety performance to PPO-Lag and significantly outperforms the PPO in terms of cost violations. Video clips showcasing the learned policies in the $3$ scenarios are included in the supplementary material.
@@ 5.3 Cost Transfer
We also demonstrate the potential of using the inferred cost to train a different agent to solve the same task from scratch and without additional feedback. We take the inferred cost function obtained during the training of a Point agent to solve the Circle and Goal tasks to train the Doggo agent. The observation space in the environments consists of agent-related telemetry (acceleration ,velocity) and task-related information (lidar sensors for goal/boundary/obstacle detection). In these experiments we do not use the agent-related information when learning the cost function, so that it can be transferred to a new agent. This cost function is kept fixed when training the second agent and the task-related features are used for cost-inference. We compare the performance of this approach with that of the PPO-Lagrangian algorithm that utilizes the underlying cost function of the environment. From Table 2 we can infer that agents trained using transferred cost function are comparable in performance to agents trained using the unknown underlying cost function.
Table 2: Comparison of the PPOLag when trained with the underlying task costs vs transferred costs. The results are averaged over 3 independent seeds.
@@
5.4 Ablation Study
Finally, we showcase the effectiveness of the proposed novelty based sampling mechanism. For this, we compare it against other querying mechanisms popular in the literature [16, 18]. These methods require a predefined budget on the number of trajectories to be shown to the evaluator. We compare against two schedules: (1) Uniform schedule: fixed number of trajectories shown for feedback after each round and (2) Decreasing schedule: number of trajectories shown decrease in proportion to $\frac{1}{t}$ each round. For each schedule, we test the following strategies for sampling the subset of trajectories to be shown to the evaluator: (a) Random sampling: sample a random subset of trajectories and (b) Entropy sampling: Trajectories are sampled in descending order of average entropy $\mathcal{H}(p^{safe}_{\theta}(s_{t},a_{t}))$ of the estimator. From Figure 2, we can see that entropy based sampling outperforms random sampling, and a decreasing schedule outperforms a uniform schedule. However, all the methods fall short of the proposed novelty based sampling mechanism.
Figure 2: Comparison of different sampling and scheduling schemes. The results are averaged over $3$ independent seeds. The proposed sampling method generates on average $1950$ queries, hence for fair comparison the other methods were given a budget of $2000$ queries.
@@ 6 Limitations, Future Work and Broader Impact
Our method relies on state-level feedback in certain environments which can be costly if obtained from humans. While our theoretical results characterize the bias induced by collecting feedback for longer horizons, we do not explore explicit solutions to mitigate it. Investigating such solutions could enhance the performance of our method and is an interesting direction for future research. The proposed research has significant societal impacts. It can enable safer autonomous systems, potentially reducing accidents and enhancing road safety. However, reliance on learned cost functions can raise ethical concerns about accountability and liability.
7 Related Work
The Constrained Markov Decision Process (CMDP) framework [4] has emerged as a valuable tool in safety-critical settings, offering a clear separation between task-related information (rewards) and safety-related information (costs). This facilitates easier environment specification and transferability of the cost function across similar environments. While previous studies have extensively explored this problem [1, 33, 32, 14], they work under the assumption of a known cost function, which can pose limitations when designing the cost function is complex or its evaluation is expensive. Therefore, there arises a need to infer the cost function from external data sources in these cases.
In the past, cost inference has mainly been done from two types of data: 1) evaluator feedback and 2) constraint abiding expert demonstrations. Prior work in learning from feedback make limiting assumptions on the problem such as deterministic transitions and smoothness [26, 30, 10] or linearity assumptions [5] on the cost function. Also, none of these works consider the problem of actively querying the human in order to minimize the feedback required. This has been subsequently explored from a theoretical perspective in [7] but the proposed algorithm has yet to be adapted for complex environments. Also, these methods are limited to feedback collected for individual state action pairs which can be costly when working with human evaluators. We address these issues with our RLSF approach by intentionally reducing feedback required and making no assumptions on the cost function.
Inferring the cost from constraint satisfying expert demonstrations has received much attention recently [6, 31, 20]. However, these expert demonstrations can be difficult to acquire in many settings. Interventions are another type of feedback that can be collected from human evaluator. [21] rely on the human to terminate the episode when the agent is about to engage in unsafe behaviour while [23] rely on the human to take over control of the agent in such a scenario and lead it to a safe state. This work is also related to the analogous problem of reward inference in regular MDPs, which use similar types of data: human feedback [11, 16, 18] and expert demonstrations [13, 12].
Another related area is that of exploration: [8] introduced a pseudo-prediction task with the idea that states with high prediction errors indicate low visitation frequency due to epistemic uncertainty. [27] calculate state-density using Simhash [9] to provide an exploration bonus for states with low visitation frequency.

### v2

@@ 4.4 Policy Optimization
After the data collection round, where we sample trajectories $\{\tau\}$ and their corresponding rewards $\{r(\tau)\}$ using $\pi$, we estimate their costs $c(\tau)$ using the inferred cost function. Our proposed method allows for the policy to be updated utilizing any on-policy constrained RL algorithm. In this study, we employ the PPO-Lagrangian algorithm[3] that combines the PPO algorithm [26] with a lagrangian multiplier to ensure safety.
@@
1:Input: cost threshold $c_{max}$, segment length $k$, novelty criterion $e$
@@
11:              for each segment $\tau_{j:j+k-1}\in\tau^{i}$ do
12:                  Obtain feedback $y^{safe}$ for the segment $\tau^{i}_{j:j+k-1}$.
13:                  $D\leftarrow D\cup\{((s,a),y^{safe})\ \hskip 5.0pt\forall(s,a)\in\tau^{i}_{j:j+k-1}\}$
14:                  $d(\phi(s))\leftarrow d(\phi(s))+1\hskip 5.0pt\forall s\in\tau^{i}_{j:j+k-1}$triangleright Update the densities
@@
22:     Infer costs $\{c_{\theta}(\tau_{i})\}$ for all $\tau^{i}\sim\{\tau\}$.
@@ 5 Experiments
We investigate the following questions in our experiments:
  1. 1.
    Does RLSF succeed in effectively learning safe behaviours?
  2. 2.
    Can the inferred cost function be transferred across agents in the same task?
  3. 3.
    How does the proposed novelty based sampling scheme compare with other methods used in the literature?
  4. 4.
    How accurate is inferred cost function compared to the true cost function?
  5. 5.
    How can we address the overestimation bias of the inferred cost function as described in Section 4.2?
@@ 5.1 Experiment Setup
Table 1: Performance of different algorithms on the Safety Benchmarks. The first $7$ environments represent the hard constraint case. The remaining environments illustrate the soft constraint case, with values in brackets indicating the cost threshold. Each algorithm is run for $6$ independent seeds. (orange) and (blue) indicate the best performance in the known costs and inferred costs settings, respectively. Algorithms with a cost violation (C.V) rate below $1\%$ are deemed to have equal performance in terms of safety.
@@
Point Circle Return 45.26 $46.09$ $36.20\pm 3.95$ $22.26\pm 9.59$ $\color[rgb]{0.04,1,1}{\textbf{36.42}}\pm\textbf{1.78}$
@@
Car Circle Return 14.34 $15.21$ $5.18\pm 2.48$ $6.34\pm 2.87$ $\color[rgb]{0.04,1,1}{\textbf{9.37}}\pm\textbf{0.97}$
@@
Biased Pendulum Return 717.43 $983.27$ $495.58\pm 160.84$ $577.15\pm 184.31$ $\color[rgb]{0.04,1,1}{\textbf{721.48}}\pm\textbf{111.49}$
@@
Blocked Swimmer Return $22.62$ 21.05 $86.96\pm 10.69$ $2.15\pm 8.58$ $\color[rgb]{0.04,1,1}{\textbf{16.09}}\pm\textbf{1.44}$
@@
HalfCheetah Return $2786.71$ 2497.82 $3031.7\pm 336.48$ $257.34\pm 147.35$ $\color[rgb]{0.04,1,1}{\textbf{2112.63}}\pm\textbf{161.26}$
@@
Hopper Return $1705.00$ 1555.25 $1097.57\pm 56.35$ $990.08\pm 8.66$ $\color[rgb]{0.04,1,1}{\textbf{1408.71}}\pm\textbf{27.3}$
@@
Walker2d Return $2947.25$ 2925.23 $2195.94\pm 134.21$ $993.38\pm 17.69$ $\color[rgb]{0.04,1,1}{\textbf{2783.29}}\pm\textbf{57.51}$
@@
Point Goal Return 26.16 $26.1$ $1.61\pm 1.8149$ $10.86\pm 4.1$ $\color[rgb]{0.04,1,1}{\textbf{24.65}}\pm\textbf{0.59}$
Cost $(40.0)$ 34.19 $31.83$ $30.57\pm 13.29$ $52.76\pm 12.85$ $\color[rgb]{0.04,1,1}{\textbf{35.08}}\pm\textbf{1.08}$
Car Goal Return $27.37$ 26.44 $1.05\pm 2.83$ $\color[rgb]{0.04,1,1}{\textbf{10.88}}\pm\textbf{7.1}$ $24.28\pm 2.1$
Cost $(40.0)$ $41.67$ 35.41 $34.71\pm 9.87$ $\color[rgb]{0.04,1,1}{\textbf{33.33}}\pm\textbf{11.26}$ $41.25\pm 2.27$
Point Push Return $6.00$ 10.84 $0.16\pm 0.14$ $3.63\pm 1.77$ $\color[rgb]{0.04,1,1}{\textbf{2.68}}\pm\textbf{1.03}$
Cost $(35.0)$ $26.08$ 26.96 $22.89\pm 5.95$ $45.43\pm 3.86$ $\color[rgb]{0.04,1,1}{\textbf{30.51}}\pm\textbf{3.4}$
Car Push Return 3.07 $2.68$ $-3.04\pm 3.3$ $1.56\pm 0.46$ $\color[rgb]{0.04,1,1}{\textbf{1.54}}\pm\textbf{0.51}$
Cost $(35.0)$ 20.53 $20.95$ $23.25\pm 7.78$ $36.55\pm 1.48$ $\color[rgb]{0.04,1,1}{\textbf{27.69}}\pm\textbf{1.19}$
@@
We evaluate RLSF on multiple continuous control benchmarks in the Safety Gymnasium environment [17] and Mujoco [30] based environments introduced in [22]. The Circle, Blocked Swimmer and Biased Pendulum environments constrain the position of the agent whereas the Half Cheetah, Hopper and Walker-2d environments constrain the velocity of the agent. The Goal and Push tasks are the most challenging as they contain static and dynamic obstacles that the agent must avoid while completing the task.All of the above environments reflect safety challenges that an agent could potentially face in real-world scenarios. Additionally, we conduct experiments using the Driver simulator introduced in [21], which presents two scenarios that an autonomous driving agent is likely to encounter on the highway: lane changes and blocked paths. In this setup, the cost function is based on multiple variables, including speed, position, and distance to other vehicles. Additionally, we introduce a third scenario: overtaking on a two-lane highway. A detailed description of the tasks can be found in the Appendix B.
@@
We split the environments into two settings: (1) hard constraint setting ($c_{max}=0$) where the safety of the policy is measured in terms of the cost violation (CV) rate defined as the number of cost violations divided by the length of the episode and (2) soft constraint setting where $c_{max}>0$. We utilize an automated script that leverages the underlying cost function to simulate the feedback provided by the evaluator.
@@
We compare the performance of our algorithm against the following baselines: Self Imitation Safe Reinforcement Learning (SIM) [16]: SIM is a state-of-the-art method in constrained RL that also supports the case where feedback is elicited from an external evaluator. Similar to RLSF, the method consists of two stages, a data collection/feedback stage and a policy optimization stage. In the first stage, a trajectory $\tau$ is labelled as good if $[r(\tau)\geq r_{good}\hskip 5.0pt\cap\hskip 5.0pt\mathbb{I}[c(\tau)\leq c_{max}]$, and labelled as bad if $[r(\tau)\geq r_{bad}\hskip 5.0pt\cup\hskip 5.0pt\mathbb{I}[c(\tau)\geq c_{max}]$, where $r_{good}$ and $r_{bad}$ are predefined thresholds on the reward. The information $\mathbb{I}[c(\tau)\leq c_{max}]$ is received from feedback. The idea is then to imitate the good trajectories and stay away from the bad trajectories. If $\rho^{\pi}$ is the occupancy measure of the current policy $\pi$ and $\rho^{G},\rho^{B}$ are the occupancy measures of the good and bad trajectories respectively, then, $\pi$ is optimized as,
@@
As an upper bound, we compare the performance of our algorithm to scenarios where the cost function is known: PPO-Lagrangian (PPOLag) [3] and SIM with known costs (SIMKC) [16]. The objective is not to surpass their performance, but to match it. Therefore, when reporting the results of these algorithms, we present the best-performing seed across multiple runs.
All results presented use novelty-based sampling unless stated otherwise. We use a segment length of $1$ in the Driver, Goal and Push environments. This is because the Safety Goal and Push environments contain small obstacles that the agent interacts with for very brief periods of time, hence requiring more fine-grained feedback. An example of this is present in Figure 11 in the Appendix. In the Driver environments, a randomly initialized policy was highly unsafe. Thus a long segment length would force the evaluator to label every segment unsafe, making cost inference infeasible. In the remaining environments, the segment length corresponds to the length of the episode. Details on the number of queries generated for feedback can be found in Table 4 in the Appendix. We grant the two baseline methods (with unknown costs) an advantage by providing feedback for every trajectory generated.
@@ 5.2 Cost Inference across various tasks
Table 1 presents the performance of the various algorithms on the benchmark environments. RLSF significantly outperforms the two baselines in terms of reward and safety in all of the environments 222except the Car Goal environment where RLSF and PPOLag marginally exceed the threshold. RLSF comes to within $\approx 80\%$ of the performance of the best run of PPOLag in $7/11$ environments thereby underscoring its effectiveness in learning safe policies.
Figure 1: Cost Violation rate of different algorithms in the Driver environment. Each algorithm is run for $6$ independent seeds, with the curves representing the mean and the shaded regions indicating the standard error.
@@
In the driving environments, we evaluate the performance of our algorithm against two baseline methods: a naive PPO policy that solely maximizes reward and the PPO-Lag algorithm. Figure 1 presents these results. Notably, our method demonstrates comparable safety performance to PPO-Lag while significantly outperforming PPO in terms of cost violations. This highlights the effectiveness of our approach in learning safe autonomous driving policies. Video clips showcasing the learned policies in the $3$ scenarios are included in the supplementary material.
@@ 5.3 Cost Transfer
We also showcase the potential of using the inferred cost function to train an agent with a different embodiment or morphology to solve the same task from scratch, without requiring any additional feedback. We utilize the inferred cost function obtained during the training of a Point agent to train the Doggo agent for the Circle and Goal tasks. The observation space in the environments consists of agent-related telemetry (acceleration, velocity) and task-related information (lidar sensors for goal/boundary/obstacle detection). In these experiments, we do not incorporate agent-related information when learning the cost function, ensuring that it can be transferred to a new agent. Next, the inferred cost function remains fixed during the training of the second agent, and the task-related features are utilized for cost inference. We compare the performance of this approach with that of the PPO-Lagrangian algorithm that utilizes the underlying cost function of the environment. From Table 2 we can infer that agents trained using transferred cost function are comparable in performance to agents trained using the true underlying cost function.
Table 2: Comparison of PPOLag performance when trained with the underlying task cost function versus the transferred cost function. Results are averaged over three independent seeds.
@@
Figure 2: Comparison of different sampling and scheduling schemes. Results are averaged over $3$ independent seeds. The proposed sampling method generates on average $\approx 1950$ queries, hence for fair comparison the other methods were given a budget of $2000$ queries.
5.4 Effect of Novelty Sampling
Next, we demonstrate the effectiveness of the proposed novelty-based sampling mechanism by comparing it with other popular querying methods from the literature [18, 20]. These methods require a predefined budget for the number of trajectories presented to the evaluator. We evaluate two schedules: (1) Uniform schedule, where a fixed number of trajectories is shown for feedback after each round, and (2) Decreasing schedule, where the number of trajectories decreases in proportion to $\frac{1}{t}$ each round. For each schedule, we test the following strategies for sampling the subset of trajectories to be presented to the evaluator: (a) Random sampling, which selects a random subset of trajectories, and (b) Entropy sampling, where trajectories are sampled in descending order of the average entropy $\mathcal{H}(p^{safe}_{\theta}(s_{t},a_{t}))$ of the estimator. As shown in Figure 2, entropy based sampling outperforms random sampling, and a decreasing schedule outperforms a uniform schedule. However, all the methods fall short compared the proposed novelty based sampling mechanism.
Figure 3: Comparing the inferred cost to the true cost.
5.5 Analyzing the Inferred Cost function
To further evaluate the quality of the inferred reward function, we plot it alongside the true underlying cost function of the environment as presented in Figure 3. In the Point Goal environment, where feedback is collected for each state, the inferred cost function closely aligns with the true cost given sufficient data. However, in the Point Circle environment, where feedback is collected at the trajectory level, the inferred cost function exhibits a slight overestimation bias, as discussed in Section 4.2. Nevertheless, from Figure 3, we can conclude that our method learns cost functions that strongly correlate with the true costs.
5.6 Bias Correction
In Section 4.2, we established that when feedback is collected over segments, the inferred cost function tends to exhibit an overestimation bias. As a result, optimizing for a policy that is deemed safe with respect to $c_{max}$ using the inferred cost function can lead to overly conservative policies. If the bias $b$, could be calculated apriori, then adjusting the safety threshold to $c_{max}+b$ would still satisfy the safety guarantees outlined in Corollary 1. However, since estimating $b$ in advance is infeasible, we propose a heuristic approach where a constant $\delta\in\mathbb{R^{+}}$ is added to $c_{max}$ to account for the bias. It’s important to note that Corollary 1 only holds if $\delta\leq b$, meaning this heuristic doesn’t guarantee that the resulting policy will always be safe. Nevertheless, we tested this approach with different $\delta$ values in the Car Circle environment, as this setting exhibited a higher overestimation bias (as reflected in Table 1). The results, shown in Figure 4, demonstrate that adding this bonus does improve performance (with higher $\delta$ values yielding better results), though it introduces the additional challenge of tuning $\delta$.
Figure 4: Increasing $c_{max}$ by $\delta$ to correct for the overestimation bias.
@@ 6 Limitations, Future Work and Broader Impact
In this work, we consider the problem of learning cost functions from feedback in scenarios where evaluating or defining such functions is either costly or infeasible. We present a novel framework which significantly reduces the burden on the evaluator by eliciting feedback over extended horizons, a challenge that has not been addressed in previous research. To further reduce their load, we propose a novelty-based sampling method that only presents previously unseen trajectories to the evaluator. Through experiments on multiple benchmark environments, we demonstrate the effectiveness of our framework and the proposed sampling algorithm in learning safe policies. However, there are a few limitations we would like to acknowledge. First, our method relies on state-level feedback in some environments, which can be expensive to obtain. Second, while we simulate feedback using the true cost function, real-world feedback is often noisy, especially when provided by human evaluators. It would be valuable to conduct experiments involving real human subjects to validate the approach.
By enabling safer autonomous systems, our approach has the potential to enhance safety across a range of applications, including autonomous vehicles, drones used for delivery, and industrial robots operating in environments like warehouses or manufacturing plants. These improvements could significantly reduce accidents, protect human operators, and increase the overall reliability of automated systems.
7 Acknowledgments
This research/project is supported by the National Research Foundation Singapore and DSO National Laboratories under the AI Singapore Programme (AISG Award No: AISG2-RP-2020-016) and Lee Kuan Yew Fellowship awarded to Pradeep Varakantham.